# Impact of the valence band on Rydberg excitons in cuprous oxide quantum wells

### v1

@@ II.3 Oscillator strengths
Combining Eqs. (16), (19), and (20), the desired oscillator strengths can be evaluated as
$f^{\mathrm{rel}}_{\pm}=$ $\bigg|\int_{-\frac{L}{2}}^{\frac{L}{2}}\left[\mathcal{M}_{\rho}^{\pm}(Z)+\mathcal{M}_{z}^{\pm}(Z)\right]\mathrm{d}Z\bigg|^{2}$ (21)
@@
and the complex conjugate of the coefficients given in Eq. (18). Details on the numerical evaluation of Eqs. (21) and (22) are presented in Sec. II.4.
Note that in case of degenerate eigenenergies, the oscillator strengths calculated with Eq. (21) for the individual degenerate states must be added.
@@ II.4 Numerical algorithm
$=$ $\frac{1}{\sqrt{\rho}}\sum_{i,j,k}c^{J,m_{J},m}_{ijk}\,B_{i}^{n}(\rho)\,B_{j}^{n}(z_{\mathrm{e}})\,B_{k}^{n}(z_{\mathrm{h}})$ (26)
@@
for the energy eigenvalues of the Hamiltonian (10) and the expansion coefficients $c^{J,m_{J},m}_{ijk}$ of the corresponding eigenfunctions. Here, the indices $\alpha$ and $\beta$ count both the elements of the total basis set composed of B-splines, angular momentum, and spin states in Eq. (24) and the expansion coefficients $c^{J,m_{J},m}_{ijk}$ when arranged as one-dimensional vectors.
@@
The generalized eigenvalue problem (27) is solved using ARPACK routines optimized for band matrices [51]. The band and even sparse structure of the resulting Hamiltonian matrix $H_{\alpha\beta}$ and overlap matrix $O_{\alpha\beta}$ arises naturally from the compact support of the B-spline functions. We use 15 B-spline functions of order $n=5$ for both the $z_{\mathrm{e}}$ and $z_{\mathrm{h}}$ directions, defined on equidistant grids, and 26 B-spline functions for the $\rho$ direction. For the $\rho$ coordinate, the grid nodes are distributed non-equidistantly with a decreasing density over the interval $0,200\,\mathrm{nm}]$, providing a finer resolution near the origin where both the potential and the wave function vary most rapidly, see Appendix A for more details. Because only basis states with $\Delta m_{F}=0,\pm 4$ are coupled, it is sufficient to use one of the sets in Eq. (12) for the numerical diagonalization of the Hamiltonian.
We are mainly interested in the optically active states with dominant $m=0$ and $m=1$ contributions. Convergence of these states is achieved when restricting the spin and angular momentum basis to the subset with $m_{F}=-7/2$, $1/2$, and $9/2$ given in Table 2 in the irreducible subspace $\Gamma_{7}^{\pm}$ and to the subset with $m_{F}=-5/2$, $3/2$, and $11/2$ in the subspace $\Gamma_{6}^{\pm}$. The total number $N$ of basis states and thus the dimension of the generalized eigenvalue problem (27) is given by the product of the number of B-splines for the $\rho$, $z_{\mathrm{e}}$, and $z_{\mathrm{h}}$ coordinates and the number of angular momentum and spin states $|m,J,m_{J}\rangle$, and was up to $N=26\times 15\times 15\times 18=105\,300$ in our calculations.
Figure 4: Exciton spectra in a cuprous oxide QW. In (a) the QW width is increased from $L=0$ to 1010\,nm for excitons described by the hydrogenlike part of the Hamiltonian given by the first two terms in Eq. (30). In (b)-(e) the non-diagonal coupling terms of the are successively switched on using the parameters $\lambda_{0,1,2,3}\in[0,1]$ in the Hamiltonian (30) (see text). The photoabsorption spectrum for transitions with circularly polarized light and considering the complete structure of the valence band is given in (f). Some states are labeled by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$. The colors of symbols or lines indicate the expectation value $\langle m\rangle$ of the angular momentum of the exciton state.
@@
By contrast, the derivative with respect to $\rho$ at $\rho=0$ is computed using the approximation $\partial_{\rho}\psi(\rho,z\mathord{=}0,Z)\approx\frac{1}{\rho}\psi(\rho,z\mathord{=}0,Z)$, which is extrapolated, analogously to Ref. [29], to $\rho=0$. To compute the integral in Eq. (21), the integrands are evaluated at 100 positions on a equidistant grid, and the integral is solved numerically using numpy.trapz [52]. Note that the most time consuming part is the diagonalization of the generalized eigenvalue problem (27).
@@ III Results and discussion
The complex structure of the cuprous oxide valence bands has a significant effect on excitons in the QW. To observe the shifts of energy levels due to the impact of the different terms in the Hamiltonian (10), we apply the following strategy. We start by taking into account the diagonal part of the Hamiltonian, $H_{\mathrm{diag}}$, for the hydrogenlike two-band model of excitons in a QW, and then switch on, term by term, the parts in the Hamiltonian, which are non-diagonal with respect to the angular momentum and spin quantum numbers of the basis. To this aim, we modify the Hamiltonian (10) to the form
@@
by introducing the auxiliary parameters $\lambda_{0,1,2,3}\in[0,1]$, which can be used to successively switch on the individual non-diagonal terms. Note that $H_{\Delta m=0}$ in Eq. (11b) has both, nonzero diagonal and non-diagonal terms in the angular momentum and spin basis $|m,J,m_{J}\rangle$. In Eq. (30), the nonzero diagonal terms of $H_{\Delta m=0}$ are considered in the hydrogenlike model, described by the first and second term, and the non-diagonal terms, which couple $J=1/2$ and $J=3/2$ states of the yellow and green exciton series, respectively, are switched on by increasing $\lambda_{0}$ from 0 to $1$. For an energy range of the lowest, yellow exciton series a sequence of spectra obtained by diagonalization of the Hamiltonian (30) is presented in Fig. 4. The figure allows for a detailed discussion of the impact of the valence band on excitons in a cuprous oxide QW of L=10L=10\,nm width.
@@
We go through Fig. 4 from left to right. Panel (a) shows spectra of the hydrogenlike two-band model for excitons in a QW of width $L$ in the range from $L=0$ to L=10L=10\,nm. The energies are given with respect to the first scattering threshold $E_{1,1}(L)$, given by the sum of the lowest quantum-confinement energies, see Eq. (6). Note that the QW width $L=0$ refers to the analytically solvable two-dimensional Coulomb problem [30, 31], where states with the same principal quantum number $n$ but different angular momentum quantum numbers $m$ are degenerate, as can be clearly seen in Fig. 4. The degeneracy is lifted, when $L$ is increased, in the transition region from strong to weak confinement [28]. Some of the states in panel (a) are marked by quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$, with $m$ being an exact quantum number due to the rotational symmetry of the hydrogenlike problem. The colors of the states indicate the expectation values $\langle m\rangle$ of the angular momentum operator throughout Fig. 4.
@@
In panels (b)–(f) the QW width is fixed to L=10L=10\,nm to study the impact of the complex valence band on the excitons at such a QW width. In panel (b) the non-diagonal terms of $H_{\Delta m=0}$, coupling the yellow ($J=1/2$) and green ($J=3/2$) excitons, are switched on by increasing $\lambda_{0}$ from 0 to $1$. The coupling between yellow and green excitons has a strong impact on states with angular quantum number $m=0$, and smaller effect on states with $m\neq 0$, the energies of those are moderately lowered. Note that for the states in this panel the rotational symmetry around the $z$ axis is still valid, and thus $m$ and $m_{J}$ are exact quantum numbers. As the eigenenergies do not depend on the sign of these quantum numbers, states with $m\neq 0$ are fourfold degenerate (note that $m_{F}$ is half-integer).
@@
In panels (c) and (d) of Fig. 4 the non-diagonal terms $H_{\Delta m=1}$ and $H_{\Delta m=2}$ in the Hamiltonian (30) are successively switched on. These coupling terms break the rotational symmetry around the $z$ axis in a way that $m$ is not a good quantum number any longer, but the quantum number $m_{F}=m+m_{J}$ for the $z$ component of the total angular momentum $\boldsymbol{F}=\boldsymbol{L}+\boldsymbol{J}$ still is. As an important consequence, the degeneracy of, e.g., some states with $m_{F}=1/2$ and $m_{F}=3/2$, marked by solid and open symbols, respectively, is reduced in panels (c) and (d) of Fig. 4 from fourfold to twofold degeneracy.
@@
Finally, in panel (e) the non-diagonal terms $H_{\Delta m_{F}=4}$ in the Hamiltonian (30), which couple basis states with $\Delta m=\pm 2$ and $\Delta m_{F}=\pm 4$ are switched on. In this panel, $m_{F}$ is not a good quantum number any longer, and states can only be classified as belonging to one of the four irreducible subspaces $\Gamma_{6}^{\pm}$ and $\Gamma_{7}^{\pm}$, as discussed in Sec. II.2. The eigenenergies are shifted due to the coupling between different $m_{F}$ states, however, there is no further lifting of degeneracies, i.e., all states remain twofold degenerate. The rightmost spectrum in Fig. 4(e) is the one with $\lambda_{0,1,2,3}=1$, where the full impact of the valence band as described by the Hamiltonian (10) is taken into account. The oscillator strengths corresponding to this spectrum are shown in Fig. 4(f). Note that the spectrum significantly differs from that of the hydrogenlike model shown at QW width L=10L=10\,nm in Fig. 4(a). This clearly illustrates that a precise description of exciton states in QWs requires taking into account the complex structure of the valence band to accurately determine all splittings of energy levels.
@@
Figure 5: Photoabsorption spectrum of a cuprous oxide quantum well of width L=10L=10\,nm for transitions with left-handed circularly polarized light ($\hat{\bm{e}}_{x}+i\hat{\bm{e}}_{y}$). The relative oscillator strengths $f_{+}^{\mathrm{rel}}$ of parity allowed transitions to states in the subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ are represented by vertical light blue and green bars, respectively. Parity forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{+}(Z)$ (orange lines) and $\mathcal{M}_{z}^{+}(Z)$ (red lines) given in Eqs. (22a) and (22b), respectively, together with their sums (black lines). Some states are marked by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$.
@@
For a more detailed discussion, the photoabsorption spectrum of a cuprous oxide QW of width L=10L=10\,nm is shown in Fig. 5. Excitons are excited by a laser beam oriented perpendicular to the QW plane with left-handed ($\hat{\bm{e}}_{x}+i\hat{\bm{e}}_{y}$) circularly polarized light. Note that due to the selection rules discussed in Sec. II.3.1 only states in the irreducible subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ can be excited. Their relative oscillator strengths $f_{+}^{\mathrm{rel}}$ are computed by evaluation of Eq. (21) and indicated in Fig. 5 by vertical light green and light blue bars, respectively. Parity-forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{+}(Z)$ (orange lines) and $\mathcal{M}_{z}^{+}(Z)$ (red lines) given in Eqs. (22a) and (22b), respectively. As can be clearly seen, they are even functions in $Z$ for the parity allowed and odd functions for the parity forbidden states. It should be noted that, contrary to the simple two-band model discussed in Ref. [29], both densities $\mathcal{M}_{\rho}^{+}(Z)$ and $\mathcal{M}_{z}^{+}(Z)$ are nonzero and constructively or destructively interfere to the sums shown as black lines in the insets of Fig. 5.

### v2

@@ II.3 Oscillator strengths
To obtain the oscillator strengths for the excitation of the twofold generate states in the irreducible subspaces with circularly polarized light, the contributions of both states must be added. Combining Eqs. (18), (21), and (22), and using the time-reversal symmetry between the two degenerate states discussed in Sec. II.2, the desired oscillator strengths can be evaluated as
$f_{\mathrm{rel}}$ $=\bigg|\int_{-\frac{L}{2}}^{\frac{L}{2}}\left[\mathcal{M}_{\rho}^{+}(Z)+\mathcal{M}_{z}^{+}(Z)\right]\mathrm{d}Z\bigg|^{2}$
$+\bigg|\int_{-\frac{L}{2}}^{\frac{L}{2}}\left[\mathcal{M}_{\rho}^{-}(Z)+\mathcal{M}_{z}^{-}(Z)\right]\mathrm{d}Z\bigg|^{2}$ (23)
@@
and the complex conjugate of the coefficients given in Eq. (20). Note that the oscillator strengths $f_{\mathrm{rel}}$ in Eq. (23) do not depend on the handedness of the circularly polarized light. Details on the numerical evaluation of Eqs. (23) and (24) are presented in Sec. II.4.
@@ II.4 Numerical algorithm
$=$ $\frac{1}{\sqrt{\rho}}\sum_{i,j,k}c^{m,J,m_{J}}_{ijk}\,B_{i}^{n}(\rho)\,B_{j}^{n}(z_{\mathrm{e}})\,B_{k}^{n}(z_{\mathrm{h}})$ (28)
@@
for the energy eigenvalues of the Hamiltonian (12) and the expansion coefficients $c^{m,J,m_{J}}_{ijk}$ of the corresponding eigenfunctions. Here, the indices $\alpha$ and $\beta$ count both the elements of the total basis set composed of B-splines, angular momentum, and spin states in Eq. (26) and the expansion coefficients $c^{m,J,m_{J}}_{ijk}$ when arranged as one-dimensional vectors.
@@
The generalized eigenvalue problem (29) is solved using ARPACK routines optimized for band matrices [60]. The band and sparse structure of the resulting Hamiltonian matrix $H_{\alpha\beta}$ and overlap matrix $O_{\alpha\beta}$ arises naturally from the compact support of the B-spline functions. We use 15 B-spline functions of order $n=5$ for both the $z_{\mathrm{e}}$ and $z_{\mathrm{h}}$ directions, defined on equidistant grids, and 26 B-spline functions for the $\rho$ direction. For the $\rho$ coordinate, the grid nodes are distributed non-equidistantly with a decreasing density over the interval $0,200\,\mathrm{nm}]$, providing a finer resolution near the origin where both the potential and the wave function vary most rapidly, see Appendix A for more details. Because only basis states with $\Delta m_{F}=0,\pm 4$ are coupled, it is sufficient to use one of the sets in Eq. (14) for the numerical diagonalization of the Hamiltonian.
We are mainly interested in the optically active states with dominant $m=0$ and $m=1$ contributions. Convergence of these states is achieved when restricting the spin and angular momentum basis to the subset with $m_{F}=-7/2$, $1/2$, and $9/2$ given in Table 2 in the irreducible subspace $\Gamma_{7}^{\pm}$ and to the subset with $m_{F}=-5/2$, $3/2$, and $11/2$ in the subspace $\Gamma_{6}^{\pm}$. States in the remaining subsets in Eq. (14) do not need to be computed as they have the same eigenenergies and the twofold degenerate eigenstates just differ by time-reversal symmetry. The total number $N$ of basis states and thus the dimension of the generalized eigenvalue problem (29) is given by the product of the number of B-splines for the $\rho$, $z_{\mathrm{e}}$, and $z_{\mathrm{h}}$ coordinates and the number of angular momentum and spin states $|m,J,m_{J}\rangle$, and was up to $N=26\times 15\times 15\times 18=105\,300$ in our calculations.
Figure 4: Exciton spectra in a cuprous oxide QW. In (a) the QW width is increased from $L=0$ to 1010\,nm for excitons described by the hydrogenlike part of the Hamiltonian given by the first two terms in Eq. (33). In (b)-(e) the nondiagonal coupling terms are successively switched on using the parameters $\lambda_{0,1,2,3}\in[0,1]$ in the Hamiltonian (33) (see text). The photoabsorption spectrum for transitions with circularly polarized light and considering the complete structure of the valence band is given in (f). Some states are labeled by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$. The colors of symbols or lines indicate the expectation value $\langle m\rangle$ of the angular momentum of the exciton state.
@@
By contrast, the derivative with respect to $\rho$ at $\rho=0$ is computed using the approximation $\partial_{\rho}\psi(\rho,z\mathord{=}0,Z)\approx\frac{1}{\rho}\psi(\rho,z\mathord{=}0,Z)$, which is extrapolated, analogously to Ref. [32], to $\rho=0$. In Eq. (23) the contributions of both twofold degenerate states to the relative oscillator strength $f_{\mathrm{rel}}$ are considered. The integrands are evaluated at 100 positions on an equidistant grid, and the integrals are solved numerically using numpy.trapz [61]. Note that the most time consuming part is the diagonalization of the generalized eigenvalue problem (29).
@@ III Results and discussion
The complex structure of the cuprous oxide valence bands has a significant effect on excitons in the QW. To observe the shifts of energy levels due to the impact of the different terms in the Hamiltonian (12), we apply the following strategy. We start by taking into account the diagonal part of the Hamiltonian,
$H_{\mathrm{hyd}}=H_{\mathrm{diag}}+H_{\Delta m=0}\delta_{J^{\prime}J}\,$, (32)
representing the hydrogenlike two-band model of excitons in a QW, and then switch on, term by term, the parts in the Hamiltonian, which are nondiagonal with respect to the angular momentum and spin quantum numbers of the basis. To this aim, we modify the Hamiltonian (12) to the form
@@
by introducing the auxiliary parameters $\lambda_{0,1,2,3}\in[0,1]$, which can be used to successively switch on the individual nondiagonal terms. Note that $H_{\Delta m=0}$ in Eq. (13b) has both, nonzero diagonal and nondiagonal terms in the angular momentum and spin basis $|m,J,m_{J}\rangle$. In Eq. (33), the nonzero diagonal terms of $H_{\Delta m=0}$ are considered in the hydrogenlike model, described by the first and second term, and the nondiagonal terms, which couple $J=1/2$ and $J=3/2$ states of the yellow and green exciton series, respectively, are switched on by increasing $\lambda_{0}$ from 0 to $1$.
@@
For an energy range of the lowest, yellow exciton series a sequence of spectra obtained by diagonalization of the Hamiltonian (33) is presented in Fig. 4. The figure allows for a detailed discussion of the impact of the valence band on excitons in a cuprous oxide QW of L=10L=10\,nm width. We go through Fig. 4 from left to right. Panel (a) shows spectra of the hydrogenlike two-band model, Eq. (32) for excitons in a QW of width $L$ in the range from $L=0$ to L=10L=10\,nm. The energies are given with respect to the first scattering threshold $E_{1,1}(L)$, given by the sum of the lowest quantum-confinement energies, see Eq. (8). Note that in the limit $L\to 0$ both the exciton energies $E$ and the threshold energy $E_{1,1}(L)$ diverge, however, the energy differences $E-E_{1,1}(L)$ remains finite. This limit refers to the analytically solvable two-dimensional Coulomb problem [33, 34], where states with the same principal quantum number $n$, but different angular momentum quantum numbers $m$ are degenerate, as can be clearly seen in Fig. 4. The degeneracy is lifted, when $L$ is increased, in the transition region from strong to weak confinement [31]. Some of the states in panel (a) are marked by quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$, with $m$ being an exact quantum number due to the rotational symmetry of the hydrogenlike problem. The colors of the states indicate the expectation values $\langle m\rangle$ of the angular momentum operator throughout Fig. 4.
@@
In panels (b)–(f) the QW width is fixed to L=10L=10\,nm to study the impact of the complex valence band on the excitons at such a QW width. In panel (b) the nondiagonal terms of $H_{\Delta m=0}$, coupling the yellow ($J=1/2$) and green ($J=3/2$) excitons, are switched on by increasing $\lambda_{0}$ from 0 to $1$. The coupling between yellow and green excitons has a strong impact on states with angular quantum number $m=0$, and smaller effect on states with $m\neq 0$, the energies of those are moderately lowered. Note that for the states in this panel the rotational symmetry around the $z$ axis is still valid, and thus $m$ and $m_{J}$ are exact quantum numbers. As the eigenenergies do not depend on the sign of these quantum numbers, states with $m\neq 0$ are fourfold degenerate.
@@
In panels (c) and (d) of Fig. 4 the nondiagonal terms $H_{\Delta m=1}$ and $H_{\Delta m=2}$ in the Hamiltonian (33) are successively switched on. These coupling terms break the rotational symmetry around the $z$ axis in a way that $m$ is not a good quantum number any longer, but the quantum number $m_{F}=m+m_{J}$ for the $z$ component of the total angular momentum $\boldsymbol{F}=\boldsymbol{L}+\boldsymbol{J}$ still is. As an important consequence, the degeneracy of, e.g., some states with $m_{F}=1/2$ and $m_{F}=3/2$, marked by solid and open symbols, respectively, is reduced in panels (c) and (d) of Fig. 4 from fourfold to twofold degeneracy.
@@
Finally, in panel (e) the nondiagonal terms $H_{\Delta m_{F}=4}$ in the Hamiltonian (33), which couple basis states with $\Delta m=\pm 2$ and $\Delta m_{F}=\pm 4$, are switched on. In this panel, $m_{F}$ is not a good quantum number any longer, and states can only be classified as belonging to one of the four irreducible subspaces $\Gamma_{6}^{\pm}$ and $\Gamma_{7}^{\pm}$, as discussed in Sec. II.2. The eigenenergies are shifted due to the coupling between different $m_{F}$ states, however, there is no further lifting of degeneracies, i.e., all states remain twofold degenerate. The rightmost spectrum in Fig. 4(e) is the one with $\lambda_{0,1,2,3}=1$, where the full impact of the valence band as described by the Hamiltonian (12) is taken into account. The relative oscillator strengths corresponding to this spectrum are shown in Fig. 4(f). Note that the spectrum significantly differs from that of the hydrogenlike model shown at QW width L=10L=10\,nm in Fig. 4(a). This clearly illustrates that a precise description of exciton states in QWs requires taking into account the complex structure of the valence band to accurately determine all splittings of energy levels.
@@
Figure 5: Photoabsorption spectrum of a cuprous oxide quantum well of width L=10L=10\,nm for transitions with circularly polarized light ($\hat{\bm{e}}_{x}\pm i\hat{\bm{e}}_{y}$). The relative oscillator strengths $f_{\mathrm{rel}}$ of parity allowed transitions to states in the subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ are represented by vertical light blue and green bars, respectively. Parity forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{\pm}(Z)$ (orange lines) and $\mathcal{M}_{z}^{\pm}(Z)$ (red lines) given in Eqs. (24a) and (24b), respectively, together with their sums (black lines). Some states are marked by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$.
@@
For a more detailed discussion, the photoabsorption spectrum of a cuprous oxide QW of width L=10L=10\,nm is shown in Fig. 5. Excitons are excited by a laser beam oriented perpendicular to the QW plane with left- or right-handed ($\hat{\bm{e}}_{x}\pm i\hat{\bm{e}}_{y}$) circularly polarized light. Note that due to the selection rules discussed in Sec. II.3.1 only states in the irreducible subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ can be excited. Their relative oscillator strengths $f_{\mathrm{rel}}$ are computed by evaluation of Eq. (23) and indicated in Fig. 5 by vertical light green and light blue bars, respectively. Parity-forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{\pm}(Z)$ (orange lines) and $\mathcal{M}_{z}^{\pm}(Z)$ (red lines) given in Eqs. (24a) and (24b), respectively. As can be clearly seen, they are even functions in $Z$ for the parity allowed and odd functions for the parity forbidden states. It should be noted that, contrary to the simple two-band model discussed in Ref. [32], both densities $\mathcal{M}_{\rho}^{\pm}(Z)$ and $\mathcal{M}_{z}^{\pm}(Z)$ are nonzero and constructively or destructively interfere to the sums shown as black lines in the insets of Fig. 5.
Table 3: Numerical values for selected states of the exciton spectrum at QW width L=10L=10\,nm considering the complex band structure described by the Hamiltonian (12). The energy eigenvalues $E$ are compared to those of the hydrogenlike model $H_{\mathrm{hyd}}$ given in Eq. (32). Approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$ are assigned where it was uniquely possible. All states are characterized by their irreducible representations. Oscillator strengths $f_{\mathrm{rel}}$ are given relative to the strongest line.
EE\,(meV) $E_{\mathrm{hyd}}\,$(meV) n~~n $m$ $N_{\mathrm{e}}$ $N_{\mathrm{h}}$ sym $f_{\mathrm{rel}}$
112.95986-112.95986 97.28582-97.28582 1 0 1 1 $\Gamma_{7}^{+}$ 0
65.16197-65.16197 24.64831-24.64831 2 0 1 1 $\Gamma_{7}^{+}$ 0
27.80125-27.80125 26.25192-26.25192 2 1 1 1 $\Gamma_{6}^{-}$ 0.39413
27.37152-27.37152 26.25192-26.25192 2 1 1 1 $\Gamma_{7}^{-}$ 0.00251
18.16818-18.16818 9.81564-9.81564 $\Gamma_{7}^{-}$ 0.01186
16.76044-16.76044 13.92396-13.92396 22 11 22 11 $\Gamma_{6}^{+}$ 0
16.04019-16.04019 13.92396-13.92396 22 1 2 11 $\Gamma_{7}^{+}$ 0
15.18094-15.18094 4.53941-4.53941 $\Gamma_{7}^{+}$ 0
13.54263-13.54263 11.17306-11.17306 33 1 1 11 $\Gamma_{7}^{-}$ $1.00000$
13.05286-13.05286 11.14293-11.14293 33 2 1 11 $\Gamma_{6}^{+}$ 0
12.99400-12.99400 6.48151-6.48151 44 0 1 11 $\Gamma_{7}^{+}$ 0
12.68154-12.68154 11.14293-11.14293 33 2-2 1 11 $\Gamma_{6}^{+}$ 0
12.22872-12.22872 11.17306-11.17306 33 1 1 11 $\Gamma_{6}^{-}$ $0.19371$
12.15775-12.15775 $\Gamma_{7}^{-}$ $0.66918$
7.42073-7.42073 6.46143-6.46143 44 2 1 11 $\Gamma_{6}^{+}$ 0
7.23267-7.23267 3.34728-3.34728 $\Gamma_{7}^{+}$ 0
7.16253-7.16253 6.72711-6.72711 44 3-3 1 11 $\Gamma_{7}^{-}$ $0.00012$
7.14464-7.14464 6.72711-6.72711 44 3-3 1 11 $\Gamma_{6}^{-}$ 0.00002
7.01147-7.01147 6.46143-6.46143 44 2-2 1 11 $\Gamma_{6}^{+}$ 0
6.88742-6.88742 6.08855-6.08855 44 1 1 11 $\Gamma_{7}^{-}$ $0.32152$
6.84615-6.84615 6.08855-6.08855 44 1 1 11 $\Gamma_{6}^{-}$ 0.09737
6.61281-6.61281 2.08170-2.08170 33 1 2 11 $\Gamma_{6}^{+}$ 0
5.99680-5.99680 2.08170-2.08170 33 1 2 11 $\Gamma_{7}^{+}$ 0
For convenience, selected energy eigenvalues $E$ of the Hamiltonian (12) at QW width L=10L=10\,nm corresponding to the right-most spectrum in Fig. 4(e) and the spectrum in Fig. 5 are given in Table 3. They are compared to the energies $E_{\mathrm{hyd}}$ of the hydrogenlike model, i.e., the right-most spectrum shown in Fig. 4(a). Where possible, approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$ are assigned to the states. The irreducible representations for the states calculated using the complete valence band and the relative oscillator strengths $f_{\mathrm{rel}}$ shown in Figs. 4(f) and 5 are also indicated in Table 3.
Figure 6: Comparison of the exciton spectra of the $\Gamma_{6}^{\pm}$ states in a cuprous oxide quantum well as a function of the well width $L$. Results obtained from the full Hamiltonian $H$ in Eq. (12) and the hydrogenlike model based on the Hamiltonian $H_{\mathrm{hyd}}$ given in Eq. (32) are presented by open and closed symbols, respectively. The threshold energy $E_{1,1}(L)$ has been subtracted from all energy eigenvalues.
So far, we have discussed the impact of the valence band on exciton spectra in a QW with fixed width L=10L=10\,nm. In Fig. 6 we finally compare the dependence of the energy eigenvalues of the full Hamiltonian (12) (open symbols) and of the hydrogenlike model in Eq. (32) (closed symbols) on the QW width $L$. The change of energy levels during the crossover from the 2D limit (strong confinement regime) to the bulk semiconductor (weak confinement) in the range $5\,\mathrm{nm}\leq L\leq 15\,\mathrm{nm}$ can be observed. Since the Bohr radius of the lowest calculated exciton state is about 4 nm, the QW width of 15 nm can be considered for this state as a weak confinement regime. States obtained within the simplified hydrogenlike model exhibit only moderate variations in their separation from the threshold level $E_{1,1}(L)$. By contrast, due to strong state-mixing effects, the levels calculated using the more complex model show a pronounced sensitivity to changes in the QW width, particularly in the regime of small $L$. Energy levels for L<5L<5\, nm are not shown because for such widths the Luttinger-Kohn model with the same material parameters as for the bulk is very approximate (see Sec. IV).